40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1208     Entry time: Tue Dec 30 18:51:18 2008
Author: rana,yoichi 
Type: Configuration 
Category: Electronics 
Subject: Illuminator Power Supply reset 
We noticed that none of the illuminators were working.

The switches were off on all the ports. After turning them on it still didn't work.

The +24 V Sorensen power supply which powers all of the illuminators had its OVP light on.
We turned it off, ramped the voltage to zero, turned it back on, and then went back to +24 V.

We then checked the operation of the illuminators; ETMY is still MIA.

Each of the illuminators sucks ~0.6-0.7 A when the (unlabeled) rheostat knob panel is set
to the "25" setting.

It seems pretty unwise, in the EMI sense, to be sending Amps of unshielded, high current,
switching supply outputs for 40m down the arms. This creates a huge antenna for radiating
the switching noise. I hereby assign minus 5 points to whoever designed this system.

Illuminator Upgrade:
- Use LEDs of a wavelength that the OSEMs don't see. LEDs are also cool so that the
  Suspension won't drift in alignment.

- Use 2 power supplies so that the power is balanced.

- Make is +/-12 V twisted AWG 14 wire so that the EMI is contained. Should also
  be shielded cable.
ELOG V3.1.3-